OpenStudy (anonymous):

Evaluate the following expression for a = -4 and c = 3 5a – 4c + 6(a – 5c) – 2(a + 3)

OpenStudy (anonymous):

more pain...

OpenStudy (anonymous):

yeah they are killing me here lol

OpenStudy (vishweshshrimali5):

see 5a - 4c + 6(a-5c) -2(a+3) = -20 -12 + 6(-4-15) -2(-4+3) = -20 -12 -114 +2 = -144 is this ur answer

OpenStudy (anonymous):

thanks theres 1 more left can u try it

OpenStudy (anonymous):

Evaluate the following expression for x = -8 and y = 5 5y – 3(5 – x) + 7x + 4(3x – 2y)